Check whether the series {Summation} n2x5/(n4+x3) , x belongs to [0, a] is uniformly convergent or not ,where a belongs to R


1
Expert's answer
2022-05-31T06:55:29-0400

ANSWER . The series "\sum_{n=1}^{\infty}\frac{n^{2}x^{5}}{n^{4}+x^{3}}" converges uniformly on "[0,a]" .

EXPLANATION.

Since "a>0" , then for all "x\in [0,a] , n\geq 1" the inequality

"0\leq \frac{n^{2}x^{5}}{n^{4}+x^{3}}\leq \frac{n^{2}a^{5}}{n^{4} }=\frac {a^{5} }{n^{2}}"

is true. The series "\sum_{n=1}^{\infty}\frac{ a^{5}}{n^{2} }=a^{5}\cdot \sum_{n=1}^{\infty}\frac{ 1}{n^{2} }" converges, because the series "\sum_{n=1}^{\infty}\frac{ 1}{n^{2} }" converges ( p-series with p=2). So, by the Weierstrass M test the series "\sum_{n=1}^{\infty}\frac{n^{2}x^{5}}{n^{4}+x^{3}}" converges uniformly on "[0,a]" .